PartnerObserving Log for 2007 May 09/10

Technical Observers: J. Hill, E. Diolaiti, A. DiPaola, R. Speziali
Telescope Operator: E. Lundin
Technical Support: J. Little, B. Curtis, R. Bertram
Instrument: LBC-Blue

Some more SDT-related observing during engineering time. Note that I'm paying little attention to the SDT queue. I'm only using convenient SDT targets to test various aspects of the telescope. We are testing new guiding software, repaired shutter timing, UI improvements, image quality across the field, etc.) I have followed the SDT email traffic, so the targets aren't random.

Opened at 02:10 UT Sky photometric. Light wind. T=7degC.

No sky flats because of problem with the shutter during twilight. Cycled the shutter power and then it worked fine. Could have been a loose cable?

Improvements: R. Speziali has repaired the shutter timing this afternoon by repositioning one of the magnets. Seems to be measuring the exposure time correctly.

A. DiPaola has made the OB browsing look only at *.ob files.

S. Gallozzi has sent a new version of the guiding software as of 9 May.

D. Ashby has made some technical improvements to the Az/El servo controllers.

03:17 - 03:29 Focus and active optics at the position of GRB070419a. Took 7 iterations because of oscillations in astigmatism, coma and spherical. Seeing was around 1 arcsec, so I'm not sure why it was oscillating. Used Vfastextra.ob in advance of the r' exposure. Had to manually add -400 nm Z4 after the first exposure, but then focus is holding nicely for ~60 minutes.

03:31 - Start OB grb0419a. It is taking 5 r' exposures of 200 sec each at each of 5 dither positions. Seems like more than Peter intended, so this is going to be a very deep image! Seeing starts out at about 0.8 arcsec and varies between 0.75 and 0.9 arcsec. The 5th image of dither 1 had some kind of guiding/tracking problem. The 5th image of dither 4 also had a jump of several arcsec in guiding/tracking. (It isn't obvious if these jumps were caused by the telescope or the guiding software.)

While John and Emiliano are all geared up to work on the field-dependent astigmatism problem.....these image look very uniform across the field and show little if any astigmatism. That's good, but it is also bad, because we have not (yet) changed anything about the LBC-Blue/SX collimation.

05:15 - Focus on 2nd M81 Active Optics field (2 iterations).

05:19 - M81BV - 1 arcsec images at airmass 1.4. (2 weeks since the last observation)

05:34 - UI freezes during Vfastextra at position of Lockman block 3. During the freeze it took 2 extrafocal images (I only pressed PLAY once.).

05:42 - Egami_Lockman_VLA_U_1 - Guiding fails during dither 1 although there seems to be a star/galaxy of suitable brightness. This made 1.3 arcsec images with a tail where the guiding ended. Manually added -200 nm of Z4 during dither 2 to focus the guide star. Seeing 1.2 arcsec at airmass 1.3. Guiding failed about 1/3 of the way through dither 5 with what looks like a good guide star. These failures on apparently good guide stars make me think this new version of the guide software is worse than the version used during April SDT.

06:21 - Vfastextra -- Note that this field has only 2 stars bright enough for pupil analysis, and one of those is wiped out by a bad column. This program could get better wavefront correction and focussing by selecting a nearby dedicated active optics field.

06:24 - Egami_Lockman_VLA_U2 - 1.4 arcsec U images at 1.4 airmasses. Guiding failed on dither 2 and dither 5.

Stop science observing for a while to work on the LBC-Red collimation model. We observed a dozen pointing stars plus M10 and M57 with LBC-Red and the Electrim camera. Collimation is within LBCFPIA range at all elevations. LBC-Red is giving arcsec quality images. While co-pointing is not completely solved the new DX collimation model affected the pointing as I expected. If we built a Red pointing model in this condition, we would be ready for one-eyed Red observing.

Lost half an hour when we ran into the building limit at azimuth -88. Then we relearned how to escape from Issue #66.

Last quarter moon has risen in a crystal clear sky.

09:11 - Vfastextra at NGC_5575

09:16 - NGC_5474_U - 1.1 arcsec U images at 1.29 airmasses. UI was partially stuck, and cleared itself at the end of the OB. But then started to execute the OB again until I stopped it.

Closed rear vent doors to keep moonlight off telescope.

09:31 - Vfastextra at NGC_5474

09:34 - NGC_5474_V - Seeing degraded to 1.5 arcsec V images at 1.36 airmasses, then improving back to 1.1 arcsec by the end of the dither pattern. UI is hung frequently but recovers at the end of the OB.

10:13 - NGC_5474_V_photo

Switched to the local sdtobs copy of Firefox2.

10:18 - NGC_5474_U_photo

10:23 - NGC_5585_U_photo - 1000 counts of sky background in 164 sec.

10:27 - NGC_5585_V_photo

10:34 - PG1633_UB and PG1633_Vgr

10:50 - V5sec at PG1633, el =52.8 with nominal collimation and azoffset=0

10:55 - V5sec with RY=-60 plus X=+2.7 and azoffset=0 (I think this is the direction needed to match Red pointing.)

10:58 - V5sec with RY=-60 plus X=+2.7 and azoffset=+64 (Red is presently using an azoffset of +170 arcsec.)

Close at 04:10 T=5degC, D=-4degC.

-- JohnHill - 10 May 2007
Topic revision: r6 - 21 Jul 2009, NormCushing
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback